pat has $3.80 in nickels and dimes. If there are 51 coins in all, how many are nickles?
Answer by rapaljer(4671)

Let x = number of nickels
51 - x = number of dimes

Value of coins = number of coins times value of each coin (in cents)
Value of nickels + Value of dimes = Total value of coins ($3.80 or 380 cents)
5(x) + 10(51 - x) = 380
5x + 510 - 10x = 380
-5x + 510 = 380
-5x + 510 - 510 = 380 - 510
-5x = -130
x=26 Nickels
51 - x = 25 Dimes

Check:
Total value of nickels = $1.30
Total value of dimes = $2.50
Total value of coins = $3.80